OpenStudy (anonymous):

If the sun rises at 6 AM and is directly overhead at 12 noon, estimate the time (hour and minutes) when a 34-foot tree will have a 14-foot shadow.

OpenStudy (anonymous):

At the point in time when the tree shadow is 14 feet in length, a light ray joining the top of the tree to that end of the shadow that is away from the base of the tree, makes an angle of \[\theta=\tan^{-1} {(14/34)} \] The following expression evaluates to the time day, AM, that the shadow will be 14 feet in length. \[12-{\theta \over ( \pi/2)}*6 = 10.508\] or 10:30 AM